The textbook says that the marginal rate of substitution is equal to the slope. Is it a rule or there can be cases when slope and mrs have different values. Can you provide any examples on the graph?
Answer #1

Yes, this is a rule.

Actually, the slope of an indifference curve is known as the Marginal Rate of Substitution (MRS). Any slope can be defined as the change in the quantity of one good due to change in the quantity of other good.

MRS explains the same. It tells us how much the consumer is willing to consume a good in relation to other good.

Let X and Y be the two goods.

then, MRSx,y = dY/dX = derivative of Y with respect to X

Below graph shows the MRS at moving from point A to B and from Point B to C

MRS = A x = Slope from ATB, AY = 60-40 = -2 from B+C , 44 = 40-20 = -1 20-30 toorn Dx 30-50 IC 20 30 50